# Scalable and robust regression models for continuous proportional data

### v1

@@ 2.3 Mixture of cobin regression and varying dispersion model
We say $Y$ follows a mixture of cobin distributions (micobin) with natural parameter $\theta\in\mathbb{R}$ and dispersion parameter $\psi\in(0,1)$, written as $Y\sim\mathrm{micobin}(\theta,\psi)$, if
@@
The micobin distribution is obtained by mixing the cobin distributions over the dispersion parameter $\lambda$, analogous to obtaining Student’s $t$ distribution as a scale mixture of normal distributions. See Figure 2 for comparisons with the cobin and beta distributions ($\psi=1$ corresponds to the limiting case). The mixture over the dispersion parameter preserves the mean structure and improves the robustness against outliers (Wang and Blei,, 2018). The choice of $(\lambda-1)\sim\mathrm{negbin}(2,\psi)$ leads to a property that $E(\lambda^{-1})=\psi$, which implies $\operatorname*{var}(Y)=\psi B^{\prime\prime}(\theta)$. Thus, $\psi$ plays a similar role as $\lambda^{-1}$ as a dispersion parameter, with the additional flexibility that $\psi$ can take any value between 0 and 1.
The micobin distribution is supported on the closed interval $[0,1]$ for any choice of $\psi\in(0,1)$, with positive densities at the boundaries $p_{\mathrm{micobin}}(0;\theta,\psi)=\psi^{2}\theta/(e^{\theta}-1)>0$ and $p_{\mathrm{micobin}}(1;\theta,\psi)=\psi^{2}\theta e^{\theta}/(e^{\theta}-1)>0$. Thus, it accommodates boundary data without needing to arbitrarily nudge boundary values to lie in the open interval $(0,1)$. Micobin differs from models that assign positive point mass at 0 and/or 1, which treat boundary values as qualitatively different from others; we refer to Blasco-Moreno et al., (2019) for a detailed discussion of structural and non-structural boundary values, focusing on ecological contexts.
Under the canonical link, the micobin regression model can be compactly written as
@@
which implies $g_{\mathrm{cobit}}\{E(y_{i}\mid x_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$. Micobin regression can be extended to allow the dispersion to vary systematically with covariates, which may differ from covariates in the mean regression component of the model (Smyth,, 1989; Smithson and Verkuilen,, 2006). The varying dispersion micobin regression, under the canonical link for mean and logit link for dispersion, can be written as
@@
where $\bm{d}_{i}\in\mathbb{R}^{d}$ is a dispersion covariate that may overlap with $\bm{x}_{i}$, and $\bm{\gamma}\in\mathbb{R}^{d}$ is the coefficient.
2.4 Extensions to mixed effects models with latent Gaussian structure
When data comes with multilevel, longitudinal, or spatial structure, cobin and micobin regression models can be naturally extended to mixed-effect models, and more generally to latent Gaussian models (Rue et al.,, 2009). For spatially indexed proportional data $y(s)$ and covariate $\bm{x}(s)$, we focus on spatial generalized linear mixed models (Diggle et al.,, 1998) with cobit link,
@@ 2.4 Extensions to mixed effects models with latent Gaussian structure
With a choice of response distribution such as cobin or micobin, this allows us to perform both the inference of $\bm{\beta}$ and prediction of response at new locations in a single modeling framework (Banerjee et al.,, 2014).
Inference for non-Gaussian spatial models (7) is often performed using Bayesian approaches implemented with MCMC, due to the challenges in inferring spatial random effects at many locations. However, generic sampling algorithms such as Metropolis-Hastings require careful tuning and face significant challenges when sampling high-dimensional latent parameters. In the following section, we propose a novel data augmentation scheme called Kolmogorov-Gamma augmentation, which converts cobin or micobin likelihoods into a conditionally normal likelihood and yields a simple Gibbs sampler that does not require tuning and with theoretical guarantees on rapid convergence.
@@ 3.1 Kolmogorov-Gamma distribution and integral identity
where =$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ denotes equality in distribution.
For the case where $b=1$ and $c=0$, $\mathrm{KG}(1,0)$ scaled by $\pi^{2}$ corresponds to the squared Kolmogorov (or squared Kolmogorov-Smirnov) distribution, the infinite convolution of independent exponential distributions (Andrews and Mallows,, 1974, §4). Since $\epsilon_{k}$ are gamma distributed, following a similar naming convention as Pólya-Gamma (Polson et al.,, 2013), we call $\kappa$ a Kolmogorov-Gamma random variable. The difference with the Pólya-Gamma is the term $k^{2}$ in the denominator in (8), which is $(k-0.5)^{2}$ for Pólya-Gamma. From relationships between exponential and gamma distributions, $\mathrm{KG}(b,c)$ with integer $b$ is equal in distribution to the sum of $b$ independent $\mathrm{KG}(1,c)$ variables. In Section 4, we study the density and random-variate generating scheme in depth.
@@
Proof.
$E\{\exp(-\kappa t)\}=\prod_{k=1}^{\infty}E\left\{\exp\left(-\frac{\epsilon_{k}t}{2\pi^{2}k^{2}}\right)\right\}=\prod_{k=1}^{\infty}\left(1+\frac{t}{2\pi^{2}k^{2}}\right)^{-b}=\left[\frac{(t/2)^{1/2}}{\sinh\{(t/2)^{1/2}\}}\right]^{b},$ (10)
Theorem 2 shows that the conditional distribution with density $p(\kappa\mid\eta)\propto e^{-\kappa\eta^{2}/2}{p_{\textsc{kg}}}(\kappa;b,0)$ is $\mathrm{KG}(b,\eta)$, which arises from treating the integrand of (9) as an unnormalized density of $\kappa$.
@@ 3.2 Kolmogorov-Gamma augmentation and blocked Gibbs sampler
Based on Theorem 1, consider the augmented model by introducing KG variables $\bm{\kappa}=\{\kappa_{i}\}_{i=1}^{n}$,
@@
for $i=1,\dots,n$, which reduces to original cobin regression upon marginalization $\int_{0}^{\infty}p(y_{i},\kappa_{i}\mid x_{i},\bm{\beta},\lambda)\mathrm{d}\kappa_{i}=p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$. Then, the log-likelihood from (11) conditional on $\bm{\kappa}$ and $\lambda$ becomes a quadratic form in terms of $\bm{\beta}$. In addition, the conditional distribution of $\kappa_{i}$ given $\bm{\beta}$ and $\lambda$ is $\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ according to Theorem 2. Under a Bayesian framework with a normal prior on $\bm{\beta}$, this leads to straightfoward Gibbs samplers. For micobin regression, the same augmentation strategy can be adopted by conditioning on the mixing variable $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})$.
Algorithms 1 and 2 describe blocked Gibbs samplers for cobin and micobin regression, where $X\in\mathbb{R}^{n\times p}$ is the design matrix. We assume a normal prior $\bm{\beta}\sim N_{p}(0,\Sigma_{\beta})$, some proper prior $\lambda\sim p_{\lambda}$ for cobin regression, and a beta prior $\psi\sim\mathrm{Beta}(a_{\psi},b_{\psi})$ for micobin regression. We set a large upper bound $L$ on $\lambda$, as the posterior probability of very large $\lambda$ is negligible in practice. The update steps for $\lambda$ and $\bm{\kappa}$ are blocked together, which improves mixing and avoids the evaluation of the density of KG in (11) when updating $\lambda$. The beta prior for $\psi$ in micobin regression leads to a conditionally conjugate update. In Appendix S.4, we describe the detailed derivations and the application of the Kolmogorov-Gamma augmentation to an EM algorithm for estimating the posterior mode, and discuss sampling strategies for varying dispersion micobin regression models (6) and spatial models (7).
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda_{i},\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda_{i},\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$triangleright steps 1,2 jointly updates $(\bm{\lambda},\bm{\kappa})$
@@
4:Sample $\psi$ from $(\psi\mid\bm{\lambda})\sim\mathrm{Beta}(a_{\psi}+2n,b_{\psi}-n+\sum_{i=1}^{n}\lambda_{i})$triangleright steps 3,4 jointly updates $(\bm{\beta},\psi)$
@@
Furthermore, we show that the proposed Gibbs sampler for cobin and micobin regression based on Kolmogorov-Gamma augmentation is uniformly ergodic, meaning that the Markov chain converges to the posterior geometrically fast in terms of total variation distance, uniformly at any initialization. This implies that the Markov chain mixes rapidly and guarantees the existence of central limit theorems for Monte Carlo averages of functions of $\bm{\beta}$.
@@
4 Sampling Kolmogorov-Gamma random variables
4.1 Kolmogorov-Gamma density
@@
Developing a reliable and efficient sampling method for Kolmogorov-Gamma random variables is essential for our methodology. A naive approximation approach, based on truncating the sum of gamma random variables in Definition 3, is computationally inefficient and prone to truncation errors. In this section, we introduce an efficient method for exact sampling of $\mathrm{KG}(1,c)$ random variates based on the alternating series method of Devroye, (1986). One can then sample $\mathrm{KG}(b,c)$ variables by summing $b$ independent $\mathrm{KG}(1,c)$ variates.
Proposition 2.
$a_{n}(x;c,t)=\begin{cases}\left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{L}(x),&0<x<t,\\ \left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{R}(x),&t\leq x,\end{cases}$ (12)
4.2 Sampling KG(1,c) using alternating series method
The alternating series method (Devroye,, 1986) is an effective sampling algorithm when the target density $p(x)$ is computationally expensive to evaluate but can be approximated from above and below by a sequence of envelope functions $\{S_{m}(x)\}_{m=0}^{\infty}$, satisfying $S_{0}(x)>S_{2}(x)>\dots>p(x)>\dots>S_{3}(x)>S_{1}(x)$. With such an envelope function in hand, exact sampling of $X\sim p$ can be achieved with the following steps: (1) draw $X\sim q$ from the proposal distribution $q$, (2) generate $U\sim\mathrm{Unif}(0,Mq(X))$ where $\|p/q\|_{\infty}\leq M$, (3) repeat until $U\leq S_{m}(X)$ for odd $m$ or $U>S_{m}(X)$ for even $n$, (4) accept $X$ if $m$ is odd, repeat from (1) again if $m$ is even.
Similar to Devroye, (1986) and Polson et al., (2013), our choice of envelope function for sampling $\mathrm{KG}(1,c)$ is $S_{m}(x)=\sum_{n=0}^{m}(-1)^{n}a_{n}(x;c,t)$. For this choice to be a valid envelope function, $a_{n}(x;c,t)$ must be monotonically decreasing in $n$ for any fixed $x>0$ and parameter $c$. The following Lemma 1 shows the valid range of $t$, an intersection between the range of $x$ where $a_{n}^{L}(x)$ and $a_{n}^{R}(x)$ are both monotonically decreasing in $n$.
Lemma 1.
For any fixed $x>0$ and $c\in\mathbb{R}$, the sequence $a_{n}(x;c,t)$ defined in (12) is monotonically decreasing in $n$ when $\log(2)/(3\pi^{2})<t<0.25$, where $\log(2)/(3\pi^{2})\approx 0.0234$.
In what follows, we assume that $t$ satisfies Lemma 1; the optimal choice of $t$ will be discussed soon. Since $S_{0}(x)=a_{0}(x;c,t)>{p_{\textsc{kg}}}(x;1,c)$ for all $x$, a natural choice of proposal $q$ that ensures $\|{p_{\textsc{kg}}}(\cdot;1,c)/q(\cdot)\|_{\infty}\leq M$ for some $M$ is $q(x;c,t)=M^{-1}a_{0}(x;c,t)$. Plugging in $a_{0}^{L}(x)$ and $a_{0}^{R}(x)$ in (12), the proposal distribution with density $q(x;c,t)$ is
$X\sim\begin{cases}\mathrm{GIG}(-1.5,c^{2},1/4)1(0<X<t)\quad&\text{with prob. }A^{L}(c,t)/\{A^{L}(c,t)+A^{R}(c,t)\}\\ \mathrm{Exp}(c^{2}/2+2\pi^{2})1(t\leq X)&\text{with prob. }A^{R}(c,t)/\{A^{L}(c,t)+A^{R}(c,t)\}\\ \end{cases}$ (13)
where $\mathrm{GIG}(p,a,b)1(0<X\leq t)$ is a generalized inverse Gaussian (GIG) distribution with density proportional to $x^{p-1}\exp\{-(ax+b/x)/2\}$ truncated to $(0,t)$, $\mathrm{Exp}(a)1(t\leq X)$ is an exponential distribution with rate $a$ truncated to $[t,\infty)$, $A^{L}(c,t)=\int_{0}^{t}a_{0}(x;c,t)\mathrm{d}x$, and $A^{R}(c,t)=\int_{t}^{\infty}a_{0}(x;c,t)\mathrm{d}x$. Compared to the Pólya-Gamma case (Polson et al.,, 2013), which involves inverse Gaussian and exponential distributions in its proposal, our proposal is slightly more complicated, involving GIG and exponential distributions. Peña and Jauch, (2025) obtain results on the exact evaluation of the cdf and the sampling of GIG random variables that have a half-integer parameter $p=-1.5$. We use these results in calculating $A^{L}(c,t)$ and sampling truncated GIG random variables.
Our $\mathrm{KG}(1,c)$ sampling algorithm is highly efficient. We show this by first investigating how often a proposal $X\sim q$ from (13) is accepted, based on the expected number of outer loop iterations $M=\int_{0}^{\infty}a_{0}(x;c,t)\mathrm{d}x=A^{L}(c,t)+A^{R}(c,t)$, and then by illustrating that with very high probability, only few series terms $S_{m}(x)$ need to be computed in the inner loop in order to decide whether to accept or reject $X$.
Proposition 3.
The following statements hold:
  1. 1.
  2. 2.
Compared to the Pólya-Gamma sampler implemented in the BayesLogit R package (Polson et al.,, 2013), which took $\sim 0.3$ seconds to generate 1 million samples from Pólya-Gamma(1,2), an Rcpp implementation of our Kolmogorov-Gamma sampler took $\sim 0.5$ seconds to generate 1 million samples from $\mathrm{KG}(1,2)$ in an Apple M1 CPU environment. Algorithm S.1 in Appendix S.2 describes the pseudocode and further details.

### v2

@@ 2.3 Micobin regression and varying dispersion model
$Y$ follows a mixture of cobin distributions (micobin) with natural parameter $\theta\in\mathbb{R}$ and dispersion parameter $\psi\in(0,1)$, written as $Y\sim\mathrm{micobin}(\theta,\psi)$, if
@@
The micobin distribution is obtained by mixing the cobin distributions over the dispersion parameter $\lambda$, analogous to obtaining Student’s $t$ distribution as a scale mixture of normal distributions. See Figure 2 for comparisons with the cobin and beta distributions ($\psi=1$ corresponds to the limiting case). The mixture over the dispersion parameter preserves the mean structure and improves the robustness against outliers (Wang and Blei, , 2018). The choice of $(\lambda-1)\sim\mathrm{negbin}(2,\psi)$ leads to a property that $E(\lambda^{-1})=\psi$, which implies $\operatorname*{var}(Y)=\psi B^{\prime\prime}(\theta)$. Thus, $\psi$ plays a similar role as $\lambda^{-1}$ as a dispersion parameter, with the additional flexibility that $\psi$ can take any value between 0 and 1. The micobin distribution is supported on the closed interval $[0,1]$ for any choice of $\psi\in(0,1)$, with positive densities at the boundaries $p_{\mathrm{micobin}}(0;\theta,\psi)=\psi^{2}\theta/(e^{\theta}-1)>0$ and $p_{\mathrm{micobin}}(1;\theta,\psi)=\psi^{2}\theta e^{\theta}/(e^{\theta}-1)>0$. Thus, it accommodates boundary data without needing to arbitrarily manipulate boundary values to lie in the open interval $(0,1)$.
A micobin regression model with link function $g$ is defined as $Y_{i}\mid\theta_{i},\psi\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{micobin}(\theta_{i},\psi)$ with $\theta_{i}=(B^{\prime})^{-1}\{g^{-1}(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})\}$ for $i=1,\dots,n$. In particular, using the canonical cobit link function $g_{\mathrm{cobit}}$, the micobin regression model can be compactly written as
@@
which implies $g_{\mathrm{cobit}}\{E(y_{i}\mid\bm{x}_{i})\}=\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}$.
Remark 2.
Micobin regression does not assign positive probability mass to the boundary values, and is therefore distinct from zero/one-inflated or censored models (Ospina and Ferrari, , 2012; Kubinec, , 2023; Kosmidis and Zeileis, , 2025). Unlike these two-part models, which treat boundary values as qualitatively different by introducing a latent mechanism to determine whether $y$ falls on the boundary, micobin regression employs a one-part modeling framework that directly models $E(y_{i}\mid\bm{x}_{i})$ without conditioning on $y_{i}\in(0,1)$. Although the micobin model assigns zero probability to boundary values due to its continuous formulation, it is suitable for settings where boundary values arise from measurement precision limits or rounding of data during preprocessing. In such cases, including when there is no substantial reason to assume that boundary values arise from a distinct mechanism, two-part models may be unnecessarily complex or even misleading; see Ramalho et al., (2011) for a related discussion between one-part and two-part models for continuous proportional data in the econometrics context.
Micobin regression can be extended to allow the dispersion to vary systematically with covariates, which may differ from covariates in the mean regression component of the model (Smyth, , 1989). The varying dispersion micobin regression, under the canonical link for mean and logit link for dispersion, can be written as
@@
where $\bm{d}_{i}\in\mathbb{R}^{d}$ is a dispersion covariate that may overlap with $\bm{x}_{i}$, and $\bm{\gamma}\in\mathbb{R}^{d}$ is the coefficient. The choice of the logit link for dispersion is particularly attractive in terms of posterior computation, which enables Pólya-Gamma data augmentation with the latent negative binomial model (Pillow and Scott, , 2012), see Supplementary Section S.4.3 for details.
An important distinction from the varying dispersion beta regression (Smithson and Verkuilen, , 2006) is that the formulation in (7) does not allow bimodal densities of $Y$ with spikes at both zero and one. While cobin and micobin have a more restrictive range of $\operatorname*{var}(Y)$ compared to the beta by excluding U-shaped densities, this restriction can be advantageous in many applications where the true conditional density of $Y$ given covariates is unlikely to be U-shaped. In contrast, U-shaped fitted distributions can arise in beta regression, especially in the varying dispersion model, due to lack of fit and sparse data.
2.4 Hierarchical extensions with latent Gaussian structure
When data come with multilevel, longitudinal, or spatial structure, cobin and micobin regression models can be naturally extended to mixed-effect models, and more generally to latent Gaussian models. For spatially indexed proportional data $y(s)$ and covariate $\bm{x}(s)$, we focus on spatial generalized linear mixed models (Diggle et al., , 1998) with the cobit link,
@@ 2.4 Hierarchical extensions with latent Gaussian structure
With a choice of response distribution such as cobin or micobin, this allows us to perform both inference for $\bm{\beta}$ and make predictions at new locations in a single modeling framework.
Inference for non-Gaussian spatial models (8) is often performed using Bayesian approaches implemented with MCMC, due to the challenges in inferring spatial random effects at many locations. However, generic sampling algorithms such as Metropolis-Hastings require careful tuning and face significant challenges when sampling high-dimensional latent parameters. In the following section, we propose a novel data augmentation scheme called Kolmogorov-Gamma augmentation, which converts cobin or micobin likelihoods into a conditionally normal likelihood and yields a simple Gibbs sampler that does not require tuning and with theoretical guarantees on uniform ergodicity.
@@ 3.1 Kolmogorov-Gamma distribution and integral identity
where $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ denotes equality in distribution.
For the case where $b=1$ and $c=0$, $\mathrm{KG}(1,0)$ scaled by $\pi^{2}$ corresponds to the squared Kolmogorov (or squared Kolmogorov-Smirnov) distribution, the infinite convolution of independent exponential distributions (Andrews and Mallows, , 1974, §4). Since $\epsilon_{k}$ are gamma distributed, following a similar naming convention as Pólya-Gamma (Polson et al., , 2013), we call $\kappa$ a Kolmogorov-Gamma random variable. The difference with the Pólya-Gamma is the term $k^{2}$ in the denominator in (9), which is $(k-0.5)^{2}$ for Pólya-Gamma. From relationships between exponential and gamma distributions, $\mathrm{KG}(b,c)$ with integer $b$ is equal in distribution to the sum of $b$ independent $\mathrm{KG}(1,c)$ variables.
@@
Theorem 2 shows that the conditional distribution with density $p(\kappa\mid\eta)\propto e^{-\kappa\eta^{2}/2}{p_{\textsc{kg}}}(\kappa;b,0)$ is $\mathrm{KG}(b,\eta)$, which arises from treating integrand of (10) as an unnormalized density of $\kappa$.
@@ 3.2 Kolmogorov-Gamma augmentation and blocked Gibbs sampler
Based on Theorem 1, consider the augmented model by introducing variables $\bm{\kappa}=\{\kappa_{i}\}_{i=1}^{n}$,
@@
for $i=1,\dots,n$, which reduces to original cobin regression upon marginalization $\int_{0}^{\infty}p(y_{i},\kappa_{i}\mid x_{i},\bm{\beta},\lambda)\mathrm{d}\kappa_{i}=p_{\mathrm{cobin}}(y_{i};\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta},\lambda^{-1})$. Then, the log-likelihood from (11) conditional on $\bm{\kappa}$ and $\lambda$ becomes a quadratic form in terms of $\bm{\beta}$. In addition, the conditional distribution of $\kappa_{i}$ given $\bm{\beta}$ and $\lambda$ is $\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$ according to Theorem 2. Under a Bayesian framework with a normal prior on $\bm{\beta}$, this leads to straightforward Gibbs samplers. For micobin regression, the same augmentation strategy can be adopted by conditioning on the mixing variable $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{n})$.
Algorithms 1 and 2 describe blocked Gibbs samplers for cobin and micobin regression, where $X\in\mathbb{R}^{n\times p}$ is the design matrix. We assume a normal prior $\bm{\beta}\sim N_{p}(0,\Sigma_{\beta})$, some proper prior $\lambda\sim p_{\lambda}$ for cobin regression, and a beta prior $\psi\sim\mathrm{Beta}(a_{\psi},b_{\psi})$ for micobin regression. We set a large upper bound $L$ on $\lambda$, as the posterior probability of very large $\lambda$ is negligible in practice. The update steps for $\lambda$ and $\bm{\kappa}$ are blocked together, which improves mixing and avoids the evaluation of the density of KG in (11) when updating $\lambda$. The beta prior for $\psi$ in micobin regression leads to a conditionally conjugate update. In Supplementary Section S.4, we describe the detailed derivations and the application of the Kolmogorov-Gamma augmentation to an EM algorithm for estimating the posterior mode, and discuss sampling strategies for varying dispersion micobin regression models (7) and spatial models (8).
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda,\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda,\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$ $\triangleright$ steps 1,2 jointly updates $(\lambda,\bm{\kappa})$
@@
2:Sample $\kappa_{i}$ from $(\kappa_{i}\mid\lambda_{i},\bm{\beta})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{KG}(\lambda_{i},\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})$, $i=1,\dots,n$ $\triangleright$ steps 1,2 jointly updates $(\bm{\lambda},\bm{\kappa})$
@@
4:Sample $\psi$ from $(\psi\mid\bm{\lambda})\sim\mathrm{Beta}(a_{\psi}+2n,b_{\psi}-n+\sum_{i=1}^{n}\lambda_{i})$ $\triangleright$ steps 3,4 jointly updates $(\bm{\beta},\psi)$
@@
Furthermore, we show that the proposed Gibbs sampler for cobin and micobin regression based on Kolmogorov-Gamma augmentation is uniformly ergodic, meaning that the Markov chain converges to the posterior geometrically fast in terms of total variation distance, uniformly at any initialization. This guarantees the existence of central limit theorems for Monte Carlo averages of functions of $\bm{\beta}$.
@@
3.3 Sampling Kolmogorov-Gamma random variables
@@
Developing a reliable and efficient sampling method for Kolmogorov-Gamma random variables is essential for our methodology. A naive approximation approach, based on truncating the sum of gamma random variables in Definition 3, is computationally inefficient and prone to truncation errors. We develop an efficient method for exact sampling of $\mathrm{KG}(1,c)$ random variates based on the alternating series method of Devroye, (1986). For integer $\lambda$, one can then sample $\mathrm{KG}(\lambda,c)$ variables by summing $\lambda$ independent $\mathrm{KG}(1,c)$ variates.
We outline key steps involved in developing the exact $\mathrm{KG}(1,c)$ sampler; full details are provided in Supplementary Section S.1. It begins by identifying two alternating series representations of the $\mathrm{KG}(1,c)$ density. We then find an optimal proposal distribution that minimizes the expected number of loop iterations, which is a mixture of a truncated generalized inverse Gaussian (GIG) distribution and a truncated exponential distribution. Compared to the Pólya-Gamma sampler implemented in the BayesLogit R package (Polson et al., , 2013), which took $\sim 0.3$ seconds to generate 1 million samples from Pólya-Gamma(1,2), an Rcpp implementation of our Kolmogorov-Gamma sampler took $\sim 0.5$ seconds to generate 1 million samples from $\mathrm{KG}(1,2)$ in an Apple M1 CPU environment.
The blocked Gibbs samplers described in Algorithms 1 and 2 require sampling $n$ KG random variables at each iteration, and exact sampling of a $\mathrm{KG}(\lambda,c)$ variate via convolution involves generating $\lambda$ i.i.d. $\mathrm{KG}(1,c)$ samples. Therefore, one of the key determinants of MCMC runtime other than dataset dimensions $(n,p)$ is the signal-to-noise ratio, which governs the parameter $\lambda$ (or latent $\{\lambda_{i}\}_{i=1}^{n}$ in micobin regression).